An electron moving at 8.21 × 103 m/s in a 6.76-T magnetic field experiences a magnetic force of 4.75×10−16 N. What angle does the velocity of the electron make with the magnetic field?


answer to 2 decimal places


1
Expert's answer
2021-05-27T18:53:03-0400

The magnitude of force is calculated from the relation:


"F=qvBsin\theta"

where

"q= charge\\F=force\\v=velocity\\B=magnetic\ field\ strength\\\theta=angle\ between\ direction\ of \ velocity\ and\ magnetic\ field\ strength"


So,

"F=qvBsin\theta\\\Rightarrow sin\theta=\dfrac{F}{qvB}=\dfrac{4.75\times10^{-16}}{(1.6\times10^{-19})(8.21\times10^3)(6.76)}=0.0535"


So,

"\theta=sin^{-1}(0.535)=32.34\degree"


Hence, the velocity of the electron makes an angle of "32.34\degree"with the magnetic field.